OpenStudy (anonymous):

Solve y=2xz^2 -xy for x. Answers: A.) x = 1/2z^2 B.) x = y/2z^2 C.) x = 1/ 2z^2 - 1 D.) x = y/ 2z^2 -y

OpenStudy (anonymous):

y = 2xz^2 - xy = x(2z^2 - y) x = y / (2z^2 - y) Answer: D

OpenStudy (anonymous):

Thank you!